how do i work out the volume of half a prism if the circumference is 1.2cm and the length is 350cm
Answer by rothauserc(4718)

we are asked to find 1/2 volume of a cylindrical prism
volume(V) = pi * r^2 * l, where pi is 3.1416, r is radius and l is length or height
in our example,
circumference is pi* diameter(d)
1.2 = 3.1416 * d
d =1.2 / 3.1416 = 0.38197097
r = 0.38197097 / 2 = 0.190985485 approx 0.19
half of V = (1/2) * 3.1416 * 0.19^2 * 350
half of V = 19.847058 cubic cm